OpenStudy (anonymous):

Use implicit differentiation to find an equation of the line tangent to the curve x^3+2xy+y^3=13 at the point (1,2)

OpenStudy (amistre64):

its pretty basic, what issue are you coming across?

OpenStudy (anonymous):

all i've done is find the implicit derivative. after that, i don't know what i'm supposed to do

OpenStudy (amistre64):

show me your implicit and let me see if I agree with it

OpenStudy (anonymous):

\[dy divdx=-3x ^{2}\div(2x+3y ^{2})\]

OpenStudy (amistre64):

\[x^3+2xy+y^3=13\] \[3x^2X'+2yX'+2xY'+3y^2Y'=0\] \[3x^2+2y+2xY'+3y^2Y'=0\] \[3x^2+2y+Y'(2x+3y^2)=0\] \[Y'(2x+3y^2)=-3x^2-2y\] \[Y'=\frac{-3x^2-2y}{2x+3y^2}\] this is what I get, havent tried to simplfy it tho

OpenStudy (amistre64):

this here is the equation for the slope of the line at any given point on the graph

OpenStudy (amistre64):

when you knw a point and a slope, you can form a line

OpenStudy (anonymous):

i thought that when you applied the product rule to 2xy, the 2 would go off to 0 when differentiating x, causing the first part to be 0 and just be left with 2y(dy/dx)?

OpenStudy (anonymous):

okay, so the equation we just got is the slope. all i have to do is insert it into the point slope formula then?

OpenStudy (amistre64):

well, the product rule itself doesnt change :) [fg]' = f'g + fg' regardless of what variables we choose right

OpenStudy (amistre64):

well, lets use the point given to find a numerical value for the slope, then it will be easier to use

OpenStudy (anonymous):

so y-2=(slope we found)(x-1)?

OpenStudy (amistre64):

\[Y'=\frac{-3x^2-2y}{2x+3y^2};(1,2)\] \[Y'=\frac{-3.1^2-2.2}{2.1+3.2^2}\] \[Y'=\frac{-3-4}{2+12}\] \[Y'=\frac{-7}{14}=-\frac{1}{2}\] and yes, the rest of what you got there is good for the line

OpenStudy (anonymous):

okay. is y=-1/2 the equation?

OpenStudy (amistre64):

and as a clarification, if we product rule the 2xy we get: [2xy]' = 2'xy +2x'y +2xy' = 0xy +2x' y +2x y' = 2y +2x y'

OpenStudy (amistre64):

no, the slope is -1/2

OpenStudy (anonymous):

ohhh okay, got it! i was thinking that the 2 went along with the x. I didnt realize theey were 3 seperate components! It all makes sense now. thank you!

OpenStudy (amistre64):

youre welcome, and good luck :)

OpenStudy (anonymous):

thank you! :)